OpenStudy (anonymous):

re-write in y-form and then find the point(s) of intersection 2x-5y=10 and 4(x-y)+12=2x-4

OpenStudy (anonymous):

y-form is y=mx+b

OpenStudy (anonymous):

so rewrite to be y=2/5x-2 and y=1/2x+4

OpenStudy (anonymous):

Just plug it into you calculator and find the intersection points

OpenStudy (anonymous):

oooh okay thank you